Jen is 13 years younger than Andre. The sum of their age in years is 137. What is Andre's age in years.
Answer by TimothyLamb(4379)

j = a - 13
j + a = 137
---
(a - 13) + a = 137
2a = 150
a = 75
---
Answer:
Andre is 75
Jen is 62
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php